Round to the nearest centFuture value (with changing years). Dixie Bank offers a certificate of deposit with an option to select your own investment p
Answer #1

a) Calculation of amount after 3 years:

Future value= amount*(1+rate)^no. Of years

Future value= 5000*(1+0.06)^3= 5955.08

Amount after three years is $5955.08

b) calculation of amount after five years:

Future value= 5000*(1+0.06)^5= 6691.13

Amount after 5 years is $6691.13

c) calculation of amount after 10 years:

Future value= 5000*(1+0.06)^10= 8954.24

Amount after 10 years is $8954.24

d) Calculation of amount after 20 years:

Amount= 5000*(1+0.06)^20= 16035.68

Amount after 20 years is $16035.68